OpenStudy (anonymous):

Create your own two-step equation with All Real Numbers as the solution. Show the steps to solving the equation to prove the solution All Real Numbers.

OpenStudy (anonymous):

@pratu043 Can You Help Me?

OpenStudy (anonymous):

Nice question. The equation should have exactly two steps?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

OK ... that's tricky. 3(4x + 2) = 2(6x + 3) 12x + 6 = 12x + 6 So this means there are infinite solutions.

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

welcome!